MEGHÍVÓ

2004. március 1-én, hétfőn, a Nagyteremben 14 órakor kezdődő Összintézeti Szemináriumunk programja:

PIERRE CARTIER (Párizs)

Polylogarithms and multiple zeta values: an overview

című előadása.

The Riemann zeta function and the logarithm Li_{0}(z) = − log(1 − z) can be interpolated by a function of two variables Li_{s}(z) = ∑_{n≥1} z^{n}.n^{s}. These functions have been little studied until 1980, except for the Spence-Abel function Li_{2}(z). They play now an important role in number theory, algebraic geometry and mathematical physics. A recent generalization, multiple polylogarithms, has been introduced by D. Zagier, related to previous work of V. Drinfeld. These new functions can be used as an introduction to the ambitious programme of ”Grothendieck motives”. There also exist deep connections with the theory of transcendental numbers.

MINDEN ÉRDEKLŐDŐT SZERETETTEL VÁRUNK!

2004. március 1-én, hétfőn, a Nagyteremben 14 órakor kezdődő Összintézeti Szemináriumunk programja:

PIERRE CARTIER (Párizs)

Polylogarithms and multiple zeta values: an overview

című előadása.

The Riemann zeta function and the logarithm Li_{0}(z) = − log(1 − z) can be interpolated by a function of two variables Li_{s}(z) = ∑_{n≥1} z^{n}.n^{s}. These functions have been little studied until 1980, except for the Spence-Abel function Li_{2}(z). They play now an important role in number theory, algebraic geometry and mathematical physics. A recent generalization, multiple polylogarithms, has been introduced by D. Zagier, related to previous work of V. Drinfeld. These new functions can be used as an introduction to the ambitious programme of ”Grothendieck motives”. There also exist deep connections with the theory of transcendental numbers.

MINDEN ÉRDEKLŐDŐT SZERETETTEL VÁRUNK!